### v1

@@ Abstract
One of today’s most significant and transformative technologies is the rapidly developing field of artificial intelligence (AI). Defined as a computer system that simulates human cognitive processes, AI is present in many aspects of our daily lives, from the self-driving cars on the road to the virtual assistants in our smartphones. The term ”black box” is often used to describe artificial intelligence (AI) because some AI systems are so complex and opaque. With millions of parameters and layers, these systems—deep neural networks in particular—make it difficult for humans to comprehend how they make judgments. Even if AI is capable of producing correct findings, questions of accountability, prejudice, and justice are raised by the opaqueness of its decision-making process. AI has a lot of potential, but it also comes with a lot of difficulties and moral dilemmas. In the context of explainable artificial intelligence (XAI), trust is crucial as it ensures that AI systems behave consistently, fairly, and ethically. In the present article, we explore XAI, reliable XAI, and several practical uses for reliable XAI. Once more, we go over the three main components—transparency, explainability, and trustworthiness of XAI—that we determined are pertinent in this situation. We present an overview of recent scientific studies that employ trustworthy XAI in relation to these fundamental components, as well as an analysis of how trustworthy XAI is applied in various application fields. In the end, trustworthiness is crucial for establishing and maintaining trust between humans and AI systems, facilitating the integration of AI systems into various applications and domains for the benefit of society.
@@
Artificial Intelligence(AI), Trustworthy XAI, Explainable Artificial Intelligence (XAI), Healthcare, Autonomous Vehicles
@@ 1 Introduction
Philosophers who sought to characterize human thought as the mechanical manipulation of symbols laid the foundations of modern-day artificial intelligence. These efforts resulted in the invention of the programmable digital computer [1] in the 1940s. Though it is now lost, Alan Turing may have published the first article on the subject of artificial intelligence in 1941, indicating that he was at least thinking about the concept at that time. The public was first introduced to Turing’s notion of the Turing test in his seminal work ”Computing Machinery and Intelligence” from 1950 [2]. In it, Turing cast doubt on the viability of building thinking machines. The term artificial intelligence (AI) was first used in 1950 [3] but its general adoption and use in healthcare has been hindered by many shortcomings of the original models. The introduction of deep learning in the early 2000s eliminated many of these limitations. We are entering a new era of technology where AI can be used in clinical practice through risk assessment models that improve workflow efficiency and diagnostic accuracy. AI systems can now analyze complex algorithms and learn on their own. The performance of AI systems has improved significantly in recent years. These new models expand on their capabilities to include text-image synthesis based on nearly any prompt, whereas previous systems focused primarily on generating facial images.
Figure 1: Domains throughout which artificial intelligence finds applications. [4]
The applicability and potential of artificial intelligence (AI) to transform business is already evident in the wide range of areas in which it is applied: In the field of natural language processing (NLP) [5], artificial intelligence (AI) makes it easier for computers to understand and generate human language, enabling tasks such as sentiment analysis, machine translation, and spam filtering. Furthermore, computer vision [6] enables computers to comprehend visual information, promoting developments in fields like object identification, facial recognition, and self-driving automobiles. Computers can now learn from data thanks to machine learning (ML), which has applications in fraud detection, recommendation systems, predictive analytics, and other fields. Robotics [7] is a branch of artificial intelligence that deals with the design, development, and use of machines. These machines are used in a variety of industries, including space exploration, manufacturing, healthcare [8], [9], and many more. Additionally, the incorporation of AI into business intelligence (BI) [10] signals better data gathering, processing, and display, which promotes data-driven decision-making and increases efficiency. AI offers advances in patient outcomes and medical developments in the healthcare industry [11],[12], [13] by assisting in illness diagnosis, therapy development, and tailored care. AI’s promise in education is evident in its ability to personalize instruction, engage students, and automate administrative tasks—all of which open the door to more customized learning opportunities. Through data-driven tactics, artificial intelligence (AI) in agriculture maximizes crop productivity, lowers expenses, and assures environmental sustainability. Similar to this, AI in manufacturing uses task automation and process optimization to increase productivity, efficiency, and quality. AI’s influence is felt not just in these domains but also in banking, retail, energy, transportation [14], handwriting detection [15], and government, where it is transforming operations, improving services, and reshaping global industry landscapes. Figure 1 illustrates the wide range of industries in which artificial intelligence finds widespread use. These industries include retail, security, healthcare, e-commerce, manufacturing, finance, transportation and logistics, and home furnishings. These apps depend on reasonably developed AI technologies including ML, NLP, and computer vision.
@@ 1.1 Third Wave of Artificial Intelligence (3AI)
The majority of commercial AI technology today is classified as ”narrow AI,” meaning that it consists of extremely specialized systems that excel at a limited number of clearly defined jobs and nothing else. Even the most amazing autonomous cars need a combination of limited artificial intelligence algorithms. The reliance of modern AI on enormous training data sets is another drawback. For example, a three-year-old child can recognize cats from only a few instances, while a standard machine learning algorithm would need to be given tens of thousands of cat photographs before it can recognize them with any degree of precision. The idea of ”Third Wave AI” emerges because artificial intelligence has to become more humanlike in a number of ways in order to overcome these constraints and realize its full potential.
@@
According to the Defense Advanced Research Projects Agency (DARPA) [17], third-wave AI systems will be able to understand context, apply that contextual awareness with common sense, and adapt to changing conditions. This will make it possible for AI systems and human users to connect in a more organic and intuitive way [17]. One of the few DARPA projects underway, XAI is anticipated to pave the way for ”third-wave AI systems,” in which computers comprehend the context and surroundings in which they function and gradually develop the underlying explanatory models necessary to describe occurrences in the actual world.
  • First Wave AI was centered on rules, logic, and constructed knowledge.
  • Big data, statistical learning, and probabilistic techniques were introduced by Second Wave AI.
  • The goal of third-wave AI is to develop common sense and contextual adaption skills.
According to Tractica [16], the worldwide market for artificial intelligence software is predicted to generate revenues of 118.6 billion by 2025, up from around 9.5 billion US dollars in 2018. The focus of this data is on developing AI systems that not only perform tasks accurately but also provide explanations of their decisions and actions in ways that humans can understand.
The third wave of AI refers to the evolution of AI technologies beyond traditional machine learning methods towards more advanced approaches that integrate reasoning, context awareness, and human-like understanding. This wave aims to develop AI systems capable of understanding and interacting with the world in more nuanced and sophisticated ways, often drawing inspiration from cognitive science and neuroscience.
Explainable AI (XAI), on the other hand, focuses on improving the transparency and interpretability of AI systems, especially machine learning models, so that humans can understand and trust the decisions made by these systems. XAI techniques aim to provide explanations of AI predictions and actions, enabling users, including developers, regulators, and end users, to understand the underlying considerations and factors that influence AI outcomes.
The relationship between the third wave of AI and XAI lies in their shared goal of creating AI systems that are not only capable of making accurate predictions or decisions but also capable of providing meaningful explanations for those predictions or decisions. By incorporating XAI techniques into the development of third-wave AI systems, developers can ensure that these systems are not only advanced in their capabilities but also transparent and interpretable, fostering trust and acceptance among users.
@@ 1.2 Concept of Explainable AI
Artificial Intelligence is often criticized for being hard to explain. Many opponents argue that it is hard to trust the results of an AI if one does not know how it can arrive at a certain result or conclusion. This issue becomes particularly problematic when AI-based programs and systems are unable to accomplish their goals. Developing explainability is necessary to boost public confidence in the computational execution. Figure 3 shows the AI in use today and the AI expected to be used tomorrow. If we are to hold the system accountable, we must take steps to address and minimize input inefficiencies [18].
Artificial intelligence is known to suffer from the ”black box” syndrome due to a lack of understanding of how the system works. This has ramifications for human confidentiality, arbitrary discrimination, obfuscation, and legitimacy. There are often underlying prejudices and tendencies present along with this lack of openness. By enhancing users’ comprehension of how AI-powered systems reason, XAI seeks to improve their performance. The goal of Transparent AI is to make artificial intelligence (AI) safer and more accessible than ever before [18]. Therefore, each of the TAI’s diverse features has to be considered separately, and their many facets need to be talked about.
AI’s use of machine learning (ML) techniques falls into two categories: white box and black box [19]. The results of a white box model are understandable to a subject matter expert. In contrast, a black box model is very difficult to explain and can be confusing even to domain experts [20]. The three criteria of interpretability, explainability, and transparency are adhered to by the XAI algorithm [21]. A model is transparent when the underlying principles of the machine learning model and decision-making can be understood ”when the process of extracting model parameters from training data and generating labels from test data is described and justified by the designers of the approach. ” Because communicating in a way that people can understand is called interpretability [22]. Currently, there is no universally accepted definition of the concept of explainability, but its importance is recognized. An alternative is ”a collection of interpretable features of a domain that can help generate a decision (such as classification or regression) about a given example.” If an algorithm follows these guidelines, it provides a basis for recording and validating the decision, improving the algorithm, and discovering new information.
@@
From an explainability perspective, academics are actively studying intelligent systems, which is a key topic. Formal compliance may occasionally depend on the system’s understanding. Many black box algorithms, as shown in Figure 4, present a trade-off between incredible learning performance (accuracy) and provable AI.
Figure 4: Performance of artificial intelligence vs explainability. [23]
@@ 1.3 Classification Tree of XAI
XAI techniques are divided into two categories: transparent and post-hoc methods. A transparent approach is one that represents the model’s capabilities and decision-making process in an easy-to-understand way [24]. Transparent models include Bayesian approaches, decision trees, linear regression, and fuzzy inference systems. Transparent approaches can be useful when the internal feature correlations are highly complex or linear. A comprehensive classification of different XAI methods and approaches related to different types of data is shown in Figure LABEL:Classification [lertvittaakumjorn2019human].
Figure 5: XAI categorization according on data type [24]
@@ 1.4 Definition of Transparency in Artificial Intelligence
In explainable artificial intelligence (XAI), transparency refers to the ability of an AI system to provide understandable justification for its decisions and actions [25]. In fact, transparency is one of the key components of explainable artificial intelligence (XAI). In many real-world applications, particularly those with large social implications, deciphering the reasoning behind an AI system’s choice is just as crucial as the decision itself. Assuring transparency in XAI helps prevent AI systems from being viewed as ”black boxes” and instead as instruments that enable users to get insightful and informative responses [26].
The general transparency of AI systems is further enhanced by the insights offered by XAI approaches. Users have the capacity to scrutinize the decision-making procedures, detect any partialities, and evaluate the dependability and equity of the model’s results [27]. Transparent systems are essential for guaranteeing accountability and ethical concerns in fields where the implications of AI choices might have large repercussions, such as healthcare, finance, and autonomous cars. XAI approaches enable users to find patterns, comprehend relationships, and discover any flaws or biases by providing insightful information about the inner workings of AI models [27]. Stakeholders are better equipped to make judgments, confirm that the model’s predictions are accurate, and take necessary action as a result of the improved openness.
The examination of the ethical criteria revealed a correlation between explainability and transparency and a number of other quality needs. The nine quality standards pertaining to explainability and openness are shown in Figure 6.
Figure 6: Qualitative standards of explainability and openness + Assistance; – Disagreements. [28]
@@ 1.5 Transparency Vs Explainability in AI
Transparency and explainability can be compared [29]. When an AI is transparent, the ”basic elements of data and decisions must be available for inspection during and after AI use,” according to McLarney et al. [30]. Transparency exists when a user can observe how decisions are made or have access to their data. Explainability, on the other hand, is about understanding why AI succeeds or fails and revealing how it draws on the knowledge and decision-making processes of the people it will affect. It provides a rational explanation for the AI’s actions. Users need to be able to understand what data is being collected, how the AI program processes it, and how it produces trustworthy outcomes for each individual affected. This straightforward explanation ignores the challenges we confront in simplifying ”black box” algorithms, the context that is lost, and the accuracy needed when giving consumers clear explanations. The question thus becomes, is minimal explainability preferable to nothing? [30]. Other important factors to consider include the belief that explanations can adequately account for the dynamic nature of the rich information ecosystem and the appropriateness of dealing with anomalies.
@@
Interestingly, while certain AI algorithms analyze data automatically, an increasing number of AI systems are designed to explain how their algorithms work and the reasoning behind certain decisions [17]. For example, the Bing search engine’s conversation mode offers brief explanations of how it works (Fig. 7). End users may occasionally find these explanations enough, but occasionally they may get confused about how an AI arrived at a specific conclusion or behaved in a specific way. It is impractical to expect people to become more computer literate when they are more perplexed by the explanation that is provided [17]. Rather, we need to either enhance the AI system or the justification.
@@ 1.6 Definition of Trustworthiness in Artificial Intelligence
Artificial intelligence (AI) systems that incorporate trustworthiness must take a multifaceted approach that takes organizational, ethical, and technical factors into account. Establishing criteria for assessing trustworthiness is the first step in this process. These measures should include accountability, security, privacy, openness, fairness, and ethical compliance. Foundational elements include transparent algorithms that provide intelligible justifications for AI-driven judgments and high-quality, impartial data. Strong security measures and privacy-preserving strategies protect sensitive data and fight off online attacks. Responsible AI usage is encouraged through the establishment of accountability systems and adherence to moral principles and governance structures. User-centric design, ongoing observation, and training guarantee that AI systems satisfy users’ demands while developing over time to retain their credibility. Organizations may create trustworthy, transparent, equitable, and ethical systems that inspire confidence in both users and stakeholders by incorporating these principles into all phases of the AI lifecycle.
Figure 8: The three primary elements of a reliable AI. [31]
The three elements depicted in Figure 8 - algorithmic ethics, data ethics, and practice ethics - come together to form trustworthy AI. These elements offer an abstraction level for ethical issues that is data-centric [31]. Numerous unresolved problems arise when trying to address ethical concerns with AI systems. In the research paper [31], authors describe the requirements for Trustworthy AI where the mentioned:
@@
    Human agency and oversight: AI systems must uphold the values of human freedom. AI systems must promote user agency, uphold fundamental rights, and enable human oversight to realize a democratic and equal society.
@@
    Security and technical robustness: Security and technical robustness lead to the prevention of harm. For an AI system to function reliably while minimizing harm, risks must be taken into account during development. This must cover any changes in the working environment as well as possible attacks on the system by adversaries.
@@
    Data protection and data governance: As a fundamental right that is particularly affected by the extensive data collection required by AI systems, data protection is also closely related to preventing harm. Preventing privacy harm also requires data governance that addresses the quality and integrity of the data used, its relevance, access protocols, and the ability to process it in a privacy-preserving way.
@@
    Transparency: Explainability and transparency are closely related requirements. The aim is to make all relevant aspects of an AI system transparent, including data, technology, and business models. In the era of ubiquitous computing, transparency is essential to support large-scale data collection and its benefits to consumers.
@@
    Diversity, non-discrimination, and fairness: Achieving trustworthy AI requires enabling inclusivity and diversity throughout AI systems. This is important not only to consider and involve all affected parties but also to ensure fair access and treatment. Fairness and this need go hand in hand.
@@
    Social and environmental welfare: In the spirit of justice and harm prevention, the environment and the broader community should be considered as stakeholders. Research into AI solutions to address global challenges should be promoted, making AI systems more environmentally friendly and sustainable. AI systems need to benefit everyone, including those who will come after us.
@@
    Accountability: The idea of fairness and the need for accountability go hand in hand: we need mechanisms to ensure accountability and accountability for AI systems and the results they produce, both during and after their development, application, and deployment.

### v2

@@ Abstract
Artificial Intelligence (AI) is an important part of our everyday lives. We use it in self-driving cars and smartphone assistants. People often call it a ”black box” because its complex systems, especially deep neural networks, are hard to understand. This complexity raises concerns about accountability, bias, and fairness, even though AI can be quite accurate. Explainable Artificial Intelligence (XAI) is important for building trust. It helps ensure that AI systems work reliably and ethically. This article looks at XAI and its three main parts: transparency, explainability, and trustworthiness. We will discuss why these components matter in real-life situations. We will also review recent studies that show how XAI is used in different fields. Ultimately, gaining trust in AI systems is crucial for their successful use in society.
@@
Artificial Intelligence(AI), XAI, Explainable Artificial Intelligence (XAI), Healthcare, Autonomous Vehicles
@@ 1 Introduction
The foundations of modern artificial intelligence were laid by philosophers who attempted to define human thought as the mechanical manipulation of symbols, which led to the development of the programmable digital computer [1] in the 1940s. Alan Turing may have written the first article on the topic of AI in 1941, though it is now lost, suggesting that he was at least considering the idea at that time.In his groundbreaking essay ”Computing Machinery and Intelligence” from 1950, Turing first presented the idea of the Turing test to the general public [2]. Turing questioned the feasibility of creating thinking robots in it. John McCarthy first used the term artificial intelligence (AI) in 1956 at the Dartmouth Conference[3], but the first models’ numerous flaws have prevented AI from being widely adopted and used in healthcare.
Many of these limitations were removed with the advent of deep learning in the early 2000s, and we are now entering a new era of technology where AI can be used in clinical practice through risk assessment models that increase diagnostic accuracy and workflow efficiency. Performance of AI systems has improved significantly in recent years, and these new models expand on their capabilities to include text-image synthesis based on almost any prompt, whereas previous systems primarily focused on generating facial images.
Figure 1: Applications of AI across various domains [4]
The diverse range of fields in which AI is being used already demonstrates its applicability and promise to revolutionize business: AI facilitates activities like sentiment analysis, machine translation, and spam filtering by making it easier for computers to comprehend and produce human language in the discipline of natural language processing (NLP) [5]. Additionally, computer vision [6] makes it possible for computers to understand visual data, which advances areas like facial recognition, object identification, and self-driving cars. Machine learning (ML), which has uses in fraud detection, recommendation systems, predictive analytics, and other domains, has made it possible for computers to learn from data. The design, development, and application of machines are the focus of the AI field of robotics [7].
Many industries, including manufacturing, healthcare, and space exploration, use various machines [8] [9]. Combining artificial intelligence with business intelligence (BI) [10] improves how businesses collect, process, and visualize data. This leads to better decision-making and increased productivity. In healthcare, AI helps diagnose diseases, develop treatments, and provide personalized care, which improves patient outcomes [11], [12], [13]. AI also plays a significant role in education by engaging students, customizing lessons, and automating administrative tasks, resulting in more personalized learning experiences. AI in agriculture increases agricultural output, reduces costs, and ensures environmental sustainability through data-driven strategies. In a similar vein, AI in manufacturing boosts output, efficiency, and quality through work automation and process optimization. AI is changing operations, enhancing services, and changing global industry landscapes in a number of sectors, including banking, retail, energy, transportation [14], handwriting detection [15], and government. AI is widely used in a wide number of industries, as seen in Figure 1. Retail, security, healthcare, e-commerce, manufacturing, banking, logistics and transportation, and home furnishings are some of these sectors. These applications rely on moderately advanced AI technology, such as computer vision, natural language processing, and machine learning.
The contributions of this research can be stated below:
  1. 1.
    Providing an overview of XAI that understands the significance of black box models. For fair and ethical purposes, XAI fosters trust among humans and AI.
  2. 2.
    Discussing Deep Learning based systems that will be consistent and trustworthy. Moreover, recent studies from the literature have been reviewed properly.
  3. 3.
    Providing guidelines regarding XAI that will be helpful for detecting problems from numerous domains.
  4. 4.
    The paper identifies and analyzes three key components of XAI: transparency, explainability, and trustworthiness. It details how these elements are essential for understanding and improving AI systems.
@@ 1.1 Third Wave of Artificial Intelligence (3AI)
Most current commercial AI technology is called ”narrow AI.” This means these systems are highly specialized and can only perform a few specific tasks. For example, even the best self-driving cars rely on limited AI systems. Another drawback of today’s AI is its reliance on large training data sets. A typical machine learning program needs tens of thousands of cat photos to recognize cats accurately, while a three-year-old child can do this with just a few examples. The idea of ”Third Wave AI” comes from the need for AI to become more humanlike in various ways to overcome these limitations and achieve its full potential.
@@
According to the Defense Advanced Research Projects Agency (DARPA) [17], third-wave AI systems will understand the context of situations, use common sense, and adapt to changes. This will create more natural and intuitive connections between AI systems and people [17]. One of DARPA’s active projects, called XAI, aims to develop these third-wave AI systems. These computers will learn about their environments and the contexts in which they work. They will also build explanations needed to clarify real-world events.
  • First Wave AI focused on rules, logic, and built knowledge.
  • Second Wave AI introduced big data, statistical learning, and probabilistic techniques.
  • The goal of third-wave AI is to develop common sense and the ability to adapt to different contexts.
Tractica [16] predicts that the global market for AI software will grow from about 9.5 billion US dollars in 2018 to 118.6 billion by 2025. This data aims to develop AI systems that can perform tasks accurately while providing explanations that people can understand.
The term ”third wave” refers to the advancement of AI technologies beyond traditional machine learning. This new phase focuses on creating more advanced systems that can understand context, reason, and think similarly to humans. It draws inspiration from cognitive science and neuroscience, aiming to build AI that can engage with the world in more complex and detailed ways.
XAI, or Explainable Artificial Intelligence, focuses on making AI systems, especially machine learning models, easier to understand. The goal is to help people trust the decisions these systems make. XAI techniques work to explain how AI makes predictions and why it behaves the way it does. This helps users, such as developers, regulators, and everyday users, grasp the key factors that affect AI results.
Developing AI systems that can make accurate predictions or decisions and explain why they did so is a key goal of both the third wave of AI and explainable AI. By using XAI methods in their design, developers can ensure that these advanced AI systems are not only effective but also easy to understand. This will help build user trust and acceptance.
@@ 1.2 Concept of Explainable AI
AI often struggles with what is called the ”black box” problem because users do not understand how it works. This can lead to issues like lack of trust, confusion, unfair treatment, and violations of privacy. AI systems can also have hidden biases. Explainable AI (XAI) aims to make AI systems easier to understand, helping users know how they make decisions. The goal of XAI is to make AI safer and more user-friendly. Therefore, we need to look at each part of AI individually and discuss its different aspects [18].
There are two main types of machine learning (ML) techniques: white box and black box models [19]. Experts can easily understand the results of a white box model. However, even specialists may find it hard to grasp the results of a black box model [20]. The XAI algorithm [21] follows three key principles: interpretability, explainability, and transparency. A model is transparent when it clearly explains how it gets its results from training data and how it generates labels from test data. Interpretability means being able to explain findings in a way that others understand [22]. While there’s no single accepted definition of explainability, its value is recognized. One definition describes it as a set of clear features that help make decisions, like classifying or predicting outcomes for specific cases. An algorithm that meets these standards helps document and verify decisions and improves itself based on new data.
@@
Researchers are studying intelligent systems to understand them better. This is an important topic. Sometimes, a system needs to understand its own workings to comply with rules. Many complex algorithms, shown in Figure 4, balance achieving high accuracy with being explainable.
Figure 4: Trade-off between AI model accuracy and explainability, highlighting the challenge of balancing performance with interpretability. [23]
@@ 1.3 Classification Tree of XAI
XAI techniques are divided into two categories: transparent and post-hoc methods. A transparent approach is one that represents the model’s capabilities and decision-making process in an easy-to-understand way [24]. Transparent models include Bayesian approaches, decision trees, linear regression, and fuzzy inference systems. Transparent approaches can be useful when the internal feature correlations are highly complex or linear. A comprehensive classification of different XAI methods and approaches related to different types of data is shown in Figure 5.
Figure 5: Categorization of Explainable AI (XAI) techniques based on data type, illustrating differences between transparent and posthoc approaches [24]
@@ 1.4 Definition of Transparency in Artificial Intelligence
Transparency in XAI is the capacity of an AI system to make its decisions and actions understandable through explanations [25]. Transparency is one of the most important aspects of XAI, rendering AI decisions interpretable and justified. Transparency allows decision-making processes to be closely examined by stakeholders, mitigating risks in applications with societal impact, such as healthcare and finance [26].
The explainability provided by XAI techniques also increases the overall transparency of AI systems. The users are able to examine the decision-making procedure, identify any bias, and analyze the reliability and fairness of the output of the model [27]. Transparent solutions are necessary in areas like the medical field, banking, and autonomous vehicles, where AI-decisions can have significant effects.
By providing meaningful information about the internal processes of AI models, XAI methods assist users in recognizing patterns, comprehending relationships, and revealing biases or errors [27]. Due to the heightened transparency, stakeholders can more effectively develop opinions, ensure the predictions made by the model are accurate, and take the necessary action.
The examination of ethical criteria showed a link between explainability, transparency, and other quality needs. Figure 6 displays nine quality standards related to explainability and openness. The key standards for AI transparency are marked with “O.” For example, O2 focuses on how to interpret models, O15 highlights the importance of traceability, and O5 and O12 ensure that users understand the information. By following these standards, AI models can provide clear and convincing explanations for their outcomes. Keeping these standards improves accountability and helps reduce bias in AI systems.
Figure 6: Key qualitative standards (O1-O15) related to explainability and transparency in AI systems, addressing user comprehension, interpretability, and traceability. [28]
@@ 1.5 Transparency Vs Explainability in AI
Explainability and transparency are similar concepts [29]. According to McLarney et al. [30], a transparent AI necessitates that ”Basic elements of data and decisions must be available for inspection during and after AI use.” Transparency is achieved when users have access to their data or can understand how decisions are made. On the other hand, explainability seeks to reveal the reasons for AI’s successes or failures and demonstrate how it utilizes the knowledge and judgment of those it affects. It provides a rational justification for the actions of the AI. Users must clearly know what data is collected, how the AI interprets this data, and how it produces reliable outcomes for each affected individual. This straightforward explanation overlooks the challenges we face when trying to clarify ”black box” algorithms, the context that may be omitted, and the accuracy needed when offering understandable explanations to customers. Therefore, the question arises: is having minimal explainability preferable to having none at all? [30]. Additionally, the belief that explanations can adequately address the dynamic nature of the rich information ecosystem and the appropriateness of managing anomalies are also vital factors to consider.
@@
It’s interesting to note that although certain AI algorithms evaluate data automatically, more and more AI systems are made to explain how their algorithms operate and the logic behind specific choices [17]. For instance, the conversation mode of the Bing search engine provides succinct explanations of its operation (Fig. 7). Sometimes, end users might find these explanations sufficient, but other times, they would be perplexed as to how an AI came to a particular conclusion or acted in a particular manner. When individuals are more confused by the explanation given, it is unrealistic to expect them to become more computer-literate [17]. Instead, we must improve the justification of the AI system.
@@ 1.6 Definition of Trustworthiness in Artificial Intelligence
Creating trustworthy AI systems requires a careful strategy that looks at organizational, ethical, and technical factors. The first step is to set clear standards for trustworthiness. These standards should include accountability, security, privacy, transparency, fairness, and ethical behavior. Using high-quality, unbiased data and clear algorithms that explain AI decisions is essential. Strong security measures and privacy practices protect sensitive information from cyberattacks.
It’s important to create accountability frameworks and follow ethical guidelines to ensure responsible AI use. By focusing on user needs and constantly monitoring and updating the systems, AI can stay reliable over time. Applying these principles across all stages of the AI process allows organizations to develop systems that are explainable, equitable, ethical, and robust, which fosters stakeholder and user trust.
Figure 8: The three key components of XAI: Algorithmic Ethics, Data Ethics, and Practice Ethics. [31]
The three elements illustrated in Figure 8—algorithmic ethics, data ethics, and practice ethics—intersect to create responsible AI. These elements define a data-centered way to handle ethical issues [31]. However, several open challenges still remain with respect to dealing with ethical issues in AI systems. In the work [31], authors give the vision of Trustworthy AI, which mentions that:
@@
    Human agency and oversight: AI systems have to enable human freedom. They need to facilitate user choice, safeguard fundamental rights, and enable human control. This will assist in developing an equitable and just society.
@@
    Security and technical robustness: Security and technical proficiency are crucial to prevent damage. In order to enable an AI system to operate efficiently and reduce risks, its creators ought to consider potential risks when designing it. They range from environmental alterations where the system will operate to attacks by malicious individuals.
@@
    Data protection and data governance: Privacy is a fundamental right that has been highly compromised with the vast amounts of data that artificial intelligence systems gather. It is necessary to protect individual privacy to prevent potential harm. For this purpose, robust data governance must be in place. This includes making sure that the information being utilized is precise and applicable. Furthermore, there is a necessity to establish definite rules for data access and how data must be treated while maintaining the integrity of privacy.
@@
    Transparency: Explainability and transparency are pretty much dependent on each other. The key objective is to make data, technology, and business models clear. In today’s age, which is the age of pervasive technology, transparency has become a must. It aids customers in comprehending the huge volumes of data collected and the ensuing benefits.
@@
    Fairness, diversity, and nondiscrimination: Including several voices in AI systems is vital to achieve XAI. We need to involve all individuals who may be affected to ensure equal treatment and access. Fairness and this requirement come hand in hand.
@@
    Social and environmental welfare: We have to think about the environment and community in seeking justice and doing no harm. We should finance research on AI solutions to global issues. This will make AI systems environmentally friendly and sustainable. AI is supposed to be for the good of all people, including future generations..
@@
    Accountability: Accountability and fairness are essential in the context of AI. We need to have systems of holding AI systems accountable for their actions and generated results. Accountability needs to be a constituent part of AI development, deployment, and use, both during and following the activities.
1.7 Impact of XAI on Zero Trust Architecture (ZTA)
Zero Trust Architecture (ZTA) is a security system that always checks every request, no matter where it comes from, and does not assume trust automatically. Explainable AI (XAI) helps ZTA by ensuring that decisions made by AI in security are clear and reasonable.
XAI is especially useful in identity verification, access control, and spotting unusual behavior. AI models analyze how users behave and identify any suspicious activities. By adding explainability, security analysts can better understand and confirm AI-driven security rules. This reduces false alarms and speeds up response times.
For example, AI-driven network monitoring systems that use ZTA principles can explain why a specific access attempt looks suspicious. This explanation builds trust in automated cybersecurity decisions [32].